Presentation on theme: "September 21, 2005Virgo Status – ESF Workshop1 Status of Virgo B. Mours."— Presentation transcript:
September 21, 2005Virgo Status – ESF Workshop1 Status of Virgo B. Mours
September 21, 2005Virgo Status – ESF Workshop2 Main injection problem (2004) Input Mode cleaner mirror back scattering: IMC instability when Power Recycling mirror aligned Solutions: »Sort term: Factor 10 attenuation between IMC and PR mirror – Only 0.7W injected in the ITF »Long term: add a Faraday isolator redesign the injection bench IMC length control signal Red: PR aligned Black: PR misaligned With the factor 10 attenuation
September 21, 2005Virgo Status – ESF Workshop3 Virgo Recent Progress October 2004 First lock in recycled configuration. C5 run: December 2004 Recycled and “Recombined” modes. 2005 Commissioning activities Alignment: »Drift controls using wavefront sensing now on all d.o.f., »Most loops are also closed. Automation, Control upgrades, robustness, Noise hunting, Commissioning slow down by instabilities (“jumps”) during the first semester, C6 and C7 commissioning runs. Summer shutdown preparation New “injection” (bench, PR mirror, ISYS alignment). Data analysis preparation Study C-run data, Collaborative data analysis (LIGO-Virgo, Bar-Virgo).
September 21, 2005Virgo Status – ESF Workshop4 Sensitivity evolution
September 21, 2005Virgo Status – ESF Workshop5 C6 run July 29 – August 13 Duty cycles: Locking: 89% Science mode: 86%
September 21, 2005Virgo Status – ESF Workshop6 C6 “Binaries Injections” Inject NS-NS Online Analysis [1.2-1.6] Solar Mass Injections seen on the online SNR max plots
September 21, 2005Virgo Status – ESF Workshop7 C6 “Burst Injections” Strange periodicity Inject Sine-Gausian Central frequency: 460Hz Nominal SNR: 15 Example of the third injection: Power filter algorithm Efficiency ~ 90% False alarm rate ~.3 event/sec Timing resolution: 1.3ms Centering of the North Bench quadrant!
September 21, 2005Virgo Status – ESF Workshop8 C6 noise budget
September 21, 2005Virgo Status – ESF Workshop9 Post C6 upgrades (I) Sensitivity from Aug 12 (End of C6) Sensitivity from Aug 27 Improvement due to Linear Alignment (4 loops closed instead of 2) DAC noise reduction PR z control noise reduction (Neo filter with roll-off @ 100 Hz) BS z control noise : alpha tuning New optical setup for B5 Reduction of laser frequency noise 2kHz bump due to detection tower pump (exciting external bench resonances)
September 21, 2005Virgo Status – ESF Workshop10 Post C6 upgrades (II) August 27 th (600kpc) with NEO2 September 12 th (850 kpc) without NEO2 Increase of modulation index => reduction of PR_zCorr impact Diffused light dumped on WE => no more bump ~ 100-400 Hz Modulation frequency tuning BS coil driver upgrade And LA tuning…
September 21, 2005Virgo Status – ESF Workshop11 C7 run September 14 - 19 Duty cycles: Locking: 70%; Science mode: 65%
September 21, 2005Virgo Status – ESF Workshop12 C7 Hardware injections NS-NS events Two online analysis: 1 hour of data ploted September 19 September 15 Injected events 1.2-1.6 S.M. Grid 1-3 S.M. Grid
September 21, 2005Virgo Status – ESF Workshop13 Virgo compared to LIGO
September 21, 2005Virgo Status – ESF Workshop14 Virgo-LIGO joint analysis Working group for burst and inspiral events Up to now work on simulated data : Project “1a”: Compare analysis pipelines on the same data sets. Project “2b”: Study the advantages of 3 sites for astrophysical sources »Sky location, Detection efficiency 3 talks and papers (GWDAW9 and Amaldi 6) Burst from galactic center
September 21, 2005Virgo Status – ESF Workshop15 Virgo- Bars joint analysis AURIGA, ROG Burst events and Stochastic signals Project starting with software injection 4 hours of data Plan to analysis C6 &C7
September 21, 2005Virgo Status – ESF Workshop16 Next steps Sept-Dec 05: “Injection” shutdown 2-3 months of work »New injection bench; Power should go up by a factor 10 »New recycled mirror: Better controls »Miscellaneous changes Followed by the new injection system commissioning 2006: Commissioning + data taking Alignment, controls,… A science run by the end of 2006 (coincidence with LIGO-S5) ? 2007 Data taking/Commissioning/Upgrades 2008-9: Virgo + 50W laser, New electronic, New mirrors ? (not yet decided) 2011(?): Advanced Virgo 200W laser? New beam geometry? New mirrors?...